1 mole of MnO4-  oxidises x moles of Ferric oxalate value of X is--

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
PLease state the medium
 
I am teling you in acidic medium
 
3 MnO4{-} + 24 H{+} + 5 FeC2O4 → 3 Mn{2+} + 12 H2O + 10 CO2 + 5 Fe{3+} 

(1 mol FeC2O4) x (3 mol MnO4{-} / 5 mol FeC2O4) = 0.6 mol MnO4{-}
 
hence
0.6 mol MnO4{-} = 1 mole ferric oxalate
 
hence
1 mol MnO4{-} = 1 / 0.6 = 10/6 = 1.66 mole os ferric oxalate
manel
5 years ago
as u r not mention the medium hence the que is wrong’we cant stse an aswer unless mention the medium so plz mention the medium which u r reactin the reagents